OpenStudy (anonymous):

for the sequence 2,9,16,23..., find n if a_n=65 a. 10 b. 11 c.9 d. 8 ** the n is little**

OpenStudy (amorfide):

An=a+d(n-1) where An is 65 a is the first term of the sequence n is what you wanna work out d is the difference between each number

OpenStudy (amorfide):

Give it a go and I will help you along the way if you go wrong

OpenStudy (anonymous):

2,9,16,23. 9=7+2 16=7+9 23=7+16

OpenStudy (anonymous):

Here is the sequence from n=1 to n=10 {{1, 2}, {2, 9}, {3, 16}, {4, 23}, {5, 30}, {6, 37}, {7, 44}, {8, 51}, {9, 58}, {10, 65}} What is the answer to your problem?

OpenStudy (anonymous):

i got 10. thank you! :)